OpenStudy (anonymous):

A 50-pound bag of feed is a mixture of two brands. Brand A costs $2.00 per pound and Brand B costs $1.00 per pound. The mixture is worth $1.34 per pound. How many pounds of each brand are in the mixture?

OpenStudy (anonymous):

you have to make the equations first

OpenStudy (amistre64):

a+b = 50 a(2) + b(1) = 50(1.34)

OpenStudy (amistre64):

b = 50 - a

OpenStudy (amistre64):

2a + (50-a) = 50(1.34) a + 50 = 50(1.34) a = 50(1.34) - 50

OpenStudy (anonymous):

a = 17 b = 33

OpenStudy (amistre64):

a = 50(1.34-1) a = 50(.34)

OpenStudy (amistre64):

yep...a = 17 b = 50-17 b = 33..... yep :)

OpenStudy (amistre64):

kecky; does that make sense?

OpenStudy (anonymous):

yep i also got the same answer

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

good :) youre a smart cookie :)

OpenStudy (anonymous):

can you do this step by step

OpenStudy (amistre64):

yes...

OpenStudy (anonymous):

can you use x and y

OpenStudy (amistre64):

what do we know? we know that "a" and "b" equal 50 pounds: a + b = 50 we should solve for one of these... i choose "b" You can use x and y if its easier for you; its just calling it a different name.

OpenStudy (amistre64):

lets use x and y: x + y = 50 ... now solve for "y" {or x, your choice}

OpenStudy (anonymous):

y=-x+50

OpenStudy (amistre64):

good: now we also know that the 50 pounds bag costs 1.34 for "each pound". we know that either x or y at this point have a price "for each pound" right?

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

lets apply that knowledge: x(2.00) means whatever x is, the price will be 2 dollars for every "x" pounds...makes sense? x(2) + y(1) = 50(1.34) Does this make sense to you? and if not, let me know....

OpenStudy (anonymous):

is it just like 2x+1y=50(1.34)

OpenStudy (anonymous):

i think amistre is true...

OpenStudy (amistre64):

that is exactly correct; now we substitute; or use the "value" we found for "y" in this equation to solve: remember y=-x+50 can you plug that into the equation for me?

OpenStudy (amistre64):

thanx tian :)

OpenStudy (anonymous):

x y or a b is just a matter of notations, you're free to use what you want

OpenStudy (anonymous):

yes 2x+1(-x+50)=1.34(50)

OpenStudy (amistre64):

very good "ochin hoorah-show"!! :) do you know how to solve for "x" with this equation now?

OpenStudy (anonymous):

2x-1x+50=67

OpenStudy (amistre64):

so far so good.... whats next?

OpenStudy (anonymous):

put to the xs and take 50-67

OpenStudy (amistre64):

lets try that "one step" at a time, so that we get it right and dont trip over anything ok? which one do you want to do first...doesnt matter, pick your favorite.

OpenStudy (anonymous):

2-1=1

OpenStudy (amistre64):

good; so 2x - 1x = 1x right? we know have: x + 50 = 67 whats next?

OpenStudy (anonymous):

67-50=17 x=17

OpenStudy (amistre64):

terrific!! now that we know for sure the value of "x"; it equals 17. we can figure out our first equation: x + y = 50 How do we know what value y is now?

OpenStudy (anonymous):

50-17=33 y=33

OpenStudy (amistre64):

exactly.... good job :) does it make better sense now? to double check our work; we can use these x and y values in the other equation we came up with: x(2) + y(1) = 50(1.34) 17(2) + 33(1) = 67 34 + 33 = 67 67 = 67 ... it checks out great :)

OpenStudy (anonymous):

@amistre64 are you a teacher? you explained the solution very well

OpenStudy (anonymous):

yes it does thank you!

OpenStudy (amistre64):

dind; nah, just old and senile :)